A bag contains 8 blue balls and 5 black balls .2 successive draws of 3 balls are made without replacement .find the probability that the first drawing will give 3 black balls and the second 3 blue balls.സ
1
Expert's answer
2017-03-26T05:07:07-0400
A bag contains 8 blue balls and 5 black balls. 2 successive draws of 3 balls are made without replacement. The probability that the first drawing will give 3 black balls is: P1 = 5/13*4/12*3/11 = 0.035 or 3.5%. And the probability that the second drawing will give 3 blue balls is P2 = 8/13*7/12*6/11 = 0.196 or 19.6%.